# Chinese Academy of Sciences, Beijing 100190, P. R. China

### v1

@@ 3 Density fluctuations from domain walls
After the production of domain walls via the Kibble mechanism Kibble:1976sj , the size of the wall quickly grows to the scale of Hubble horizon $r_{H}=1/H$. Therefore, the domain wall network is a few-body system, which can lead to large density fluctuations. Consider a network consisting of $N$ identical walls with mass $M_{w}$ in a Hubble volume $V_{H}\equiv 4\pi r_{H}^{3}/3$ (this assumption is reasonable at least for domain walls in the scaling regime, where domain walls have a similar dimension $\overline{L}\simeq t$). We employ the Poisson distribution to describe the probability of finding $k$ walls in a volume $V$
$\mathbb{P}(k,V)=\frac{(\bar{n}V)^{k}}{k!}e^{-(\bar{n}V)},$ (10)
where $\bar{n}=N/V_{H}$ is the average number density over the Hubble horizon volume $V_{H}$ (or the background number density). As expected, the ensemble average of the number density $n=k/V$ in a volume $V$ is equal to the background number density,
$\langle n\rangle=\sum_{k=0}^{\infty}n\mathbb{P}(k,V)=\sum_{k=0}^{\infty}\left(\frac{k}{V}\right)\mathbb{P}(k,V)=\bar{n}.$ (11)
The mean square number density fluctuations or the variance from the background is given by
@@
$\langle(\Delta n)^{2}\rangle=\langle n^{2}\rangle-\bar{n}^{2}=\sum_{k=0}^{\infty}\left(\frac{k}{V}\right)^{2}\mathbb{P}(k,V)-\bar{n}^{2}=\frac{\bar{n}}{V}.$ (12)
Let $\delta=\Delta\rho_{w}/\bar{\rho}_{\rm tot}$ be the relative energy density perturbation (density contrast), where the total background energy density $\bar{\rho}_{\rm tot}=\bar{\rho}_{w}+\bar{\rho}_{\rm rad}$, with $\bar{\rho}_{\rm rad}$ being the radiation background energy density. We then define the relative mean square energy density fluctuations (the variance of density perturbation) due to the Poisson fluctuations in the volume $V$ as
$\sqrt{\langle\delta^{2}\rangle}=\frac{\sqrt{\langle(\Delta\rho_{w})^{2}\rangle}}{\bar{\rho}_{\rm tot}}=\frac{\sqrt{\langle(\Delta\rho_{w})^{2}\rangle}}{\bar{\rho}_{w}}\frac{\bar{\rho}_{w}}{\bar{\rho}_{\rm tot}}.$ (13)
Assuming that the domain wall is the only source for the density fluctuation, i.e., $\rho_{\rm rad}=\bar{\rho}_{\rm rad}$, we have $\Delta\rho_{w}=\rho_{w}+\rho_{\rm rad}-\bar{\rho}_{\rm tot}=\rho_{w}-\bar{\rho}_{w}$. The density fluctuations can be written as
$\langle(\Delta\rho_{w})^{2}\rangle=\langle\rho_{w}^{2}\rangle-\bar{\rho}_{w}^{2}=M_{w}^{2}\langle(\Delta n)^{2}\rangle$ (14)
where $\rho_{w}=nM_{w}$. Then the relative density fluctuations of the total density in a volume $V$ are given by
$\sqrt{\langle\delta^{2}\rangle}=\frac{1}{\sqrt{\bar{n}V}}\frac{f_{w}}{f_{w}+1}\simeq\frac{f_{w}}{\sqrt{\bar{n}V}},$ (15)
where the energy density fraction of the domain wall $f_{w}=\bar{\rho}_{w}/\bar{\rho}_{\rm rad}\ll 1$ in a radiation-dominated universe.
Then the probability density distribution function of the density contrast, $P(\delta)$, is featured by a zero ensemble averaged value, $\langle\delta\rangle=0$, and a variance $\sigma$ given by Eq. (15). In this work we consider a Gaussian density distribution function
$P(\delta)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{\delta^{2}}{2\sigma^{2}}\right).$ (16)
We consider the horizon-size density perturbation induced by the domain walls in the scaling regime, the variance of the probability distribution is given by
@@
$\sigma=\left.\sqrt{\langle\delta_{w}^{2}\rangle}\right|_{V\to V_{H}}=\frac{f_{w}}{\sqrt{\bar{n}V_{H}}}=\frac{f_{w}}{\sqrt{N}}.$ (17)
@@
Simulations Press:1989yh ; Hiramatsu:2012sc ; Lazanu:2015fua indicate that the domain walls tend to merge with the neighboring ones in the evolution, resulting in a wall network with a small number of horizon-size walls. For a model with a $Z_{2}$ symmetry, there eventually exists $N=1$ domain wall that separates two domains Press:1989yh . For models with $\mathcal{N}\sim 10$ domains, one expects that $N\sim\mathcal{N}$ domain walls exist in the late stage of the network evolution. Therefore, once the domain walls contribute significantly to the total energy density, $f_{w}\sim 0.1$, the energy density perturbation due to the number density fluctuations of the domain wall can be remarkable.

### v2

@@ 3 Density fluctuations from domain walls
After the production of domain walls via the Kibble mechanism [6], the size of the walls rapidly grows to a superhorizon scale through the merger and expansion of the walls. Subsequently, the walls re-enter the horizon due to the combined effects of chopping and the pressure from the bias potential. During the horizon crossing of the domain walls, the network ultimately becomes a few-body system composed of individual domain walls. Consider a domain wall network consisting of $N$ identical walls, each with mass $M_{w}$, randomly distributed within the Hubble volume $V_{H}\equiv 4\pi r_{H}^{3}/3$, where $r_{H}=1/H$. We employ the Poisson distribution to describe the probability of finding $n$ walls within a volume of $V\equiv 4\pi r^{3}/3$
$\mathbb{P}_{n}(V)=\left(\frac{V}{V_{w}}\right)^{n}\frac{e^{-V/V_{w}}}{n!},$ (3.1)
where $V_{w}=4\pi\bar{L}^{3}/3$ is the volume of a single wall, and $V=nV_{w}$. The applicability of the Poisson distribution to the domain wall network has been validated through simulations in Refs. [51, 52]. The spatially averaged energy density of the domain wall within the volume $V$ is given by $\bar{\rho}_{w}(r)=nM_{w}/V$. The ensemble average of $\bar{\rho}_{w}(r)$ is then expressed as
$\langle\bar{\rho}_{w}(r)\rangle=\sum_{n=0}^{\infty}\bar{\rho}_{w}(r)\mathbb{P}_{n}(V)=\frac{M_{w}}{V_{w}}.$ (3.2)
Therefore, the quantity $\langle\bar{\rho}_{w}(r)\rangle\equiv\langle\bar{\rho}_{w}\rangle$ is independent of the position $r$. The density perturbation arising from Poisson fluctuations is defined as
$\delta\bar{\rho}_{w}(r)=\bar{\rho}_{w}(r)-\langle\bar{\rho}_{w}\rangle.$ (3.3)
The variance of the energy density is then given by
$\langle(\delta\bar{\rho}_{w}(r))^{2}\rangle=\langle\rho_{w}^{2}(r)\rangle-\langle\bar{\rho}_{w}\rangle^{2}=\sum_{n=0}^{\infty}\rho_{w}^{2}(r)\mathbb{P}_{n}(V)-\langle\bar{\rho}_{w}\rangle^{2}=\frac{M_{w}^{2}}{V_{w}V}.$ (3.4)
We define the relative energy density fluctuation, or the density contrast, as $\delta=\delta\bar{\rho}/\rho_{\rm tot}$. For a two-component Universe consisting of radiation and domain walls, the total density perturbation is given by $\delta\bar{\rho}=\delta\bar{\rho}_{w}+\delta\bar{\rho}_{r}$, where $\delta\bar{\rho}_{w}$ and $\delta\bar{\rho}_{r}$ are the perturbations induced by the domain wall and radiation, respectively. In this work, we neglect the radiation density perturbation, and the mean square of the density contrast is then given by
$\langle\delta^{2}(r)\rangle=\frac{\langle(\delta\bar{\rho}_{w}(r))^{2}\rangle}{\rho_{\rm tot}^{2}}=\frac{\langle\bar{\rho}_{w}\rangle^{2}V_{w}}{\rho_{\rm tot}^{2}V},$ (3.5)
where we have used Eqs. (3.2) and (3.4) for the second equality. If we consider the fluctuation within the Hubble volume, i.e., $V\to V_{H}$, then we obtain
$\left.\langle\delta^{2}\rangle\right|_{V\to V_{H}}=\frac{f_{w}^{2}}{N},$ (3.6)
where we have used the relations $f_{w}(t)=\langle\bar{\rho}_{w}\rangle/\rho_{\rm rad}\simeq\langle\bar{\rho}_{w}\rangle/\rho_{\rm tot}$ and $V_{H}=NV_{w}$.
Let us now calculate the power spectrum of the density perturbation arising from Poisson fluctuations. Denote the domain wall energy density at a location $\bm{x}$ within the volume $V$ as $\rho_{w}(\bm{x})$. The spatially averaged quantities can then be expressed as
$\bar{\rho}_{w}(r)=\frac{1}{V}\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}\rho_{w}(\bm{x})\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\delta\rho}_{w}(r)=\frac{1}{V}\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}\delta\rho_{w}(\bm{x}).$ (3.7)
Using the definition $\delta\rho_{w}(\bm{x})=\rho_{w}(\bm{x})-\langle\bar{\rho}_{w}\rangle$, we find that $\overline{\delta\rho}_{w}(r)\equiv\delta\bar{\rho}_{w}(r)=\bar{\rho}_{w}(r)-\langle\bar{\rho}_{w}\rangle$ and $\langle(\overline{\delta\rho}_{w}(r))^{2}\rangle\equiv\langle(\delta\bar{\rho}_{w}(r))^{2}\rangle=\langle\bar{\rho}_{w}^{2}(r)\rangle-\langle\bar{\rho}_{w}\rangle^{2}$. Together with Eq. (3.4), we obtain
@@
$\left\langle\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}^{\prime}\delta\rho_{w}(\bm{x})\delta\rho_{w}(\bm{x}^{\prime})\right\rangle=\frac{M_{w}^{2}V}{V_{w}}.$ (3.8)
The randomness of the spatial distribution implies that fluctuations at two different locations are independent [53]. Furthermore, the spatial distribution is assumed to be homogeneous and isotropic. With these properties, we can express the density correlator as
$\langle\delta\rho_{w}(\bm{x})\delta\rho_{w}(\bm{x}^{\prime})\rangle=g\,\delta(\bm{x}-\bm{x}^{\prime}),$ (3.9)
where $\delta(\bm{x}-\bm{x}^{\prime})$ is the Dirac delta function, and $g$ may depend on $|\bm{x}-\bm{x}^{\prime}|$. Integrating over the volume $V$, we obtain
$\int_{V}\mathrm{d}^{3}\bm{x}\int_{V}\mathrm{d}^{3}\bm{x}^{\prime}\,\langle\delta\rho_{w}(\bm{x})\delta\rho_{w}(\bm{x}^{\prime})\rangle=Vg.$ (3.10)
Using Eq. (3.8), we find $g=M_{w}^{2}/V_{w}=\langle\bar{\rho}_{w}\rangle^{2}V_{w}$. Substituting this into the correlator, we obtain the density contrast correlator as
$\left\langle\frac{\delta\rho_{w}(\bm{x})}{\rho_{\rm tot}}\frac{\delta\rho_{w}(\bm{x}^{\prime})}{\rho_{\rm tot}}\right\rangle=f_{w}^{2}V_{w}\,\delta(\bm{x}-\bm{x}^{\prime}).$ (3.11)
Expanding the density perturbation in the Fourier space, we have
$\delta_{w}(t)\equiv\frac{\delta\rho_{w}(\bm{z})}{\bar{\rho}_{w}}=\int\frac{\mathrm{d}^{3}\bm{k}}{(2\pi)^{3/2}}\,\delta_{w,\bm{k}}(t)\,e^{i\bm{k}\cdot\bm{z}},$ (3.12)
where $\bm{k}$ and $\bm{z}$ denote the wave number and coordinates in the comoving space, respectively. Using Eq. (3.11), the correlator in the comoving coordinates is given by
$\left\langle\frac{\delta\rho_{w}(\bm{z})}{\bar{\rho}_{w}}\frac{\delta\rho_{w}^{*}(\bm{z}^{\prime})}{\bar{\rho}_{w}}\right\rangle=\frac{4\pi}{3}\left(\frac{\bar{L}}{a}\right)^{3}\delta(\bm{z}-\bm{z}^{\prime}).$ (3.13)
The reduced power spectrum $\mathcal{P}_{\delta_{w}}(k)$ is defined in terms of
@@
$\left\langle\delta_{w,\bm{k}}\delta_{w,\bm{k}^{\prime}}^{*}\right\rangle\equiv\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{\delta_{w}}(k)\delta\left(\bm{k}-\bm{k}^{\prime}\right).$ (3.14)
With Eqs. (3.12) and (3.13), we obtain
$\mathcal{P}_{\delta_{w}}(k)=\frac{2}{3\pi}\left(\frac{k}{k_{e}}\right)^{3}\simeq\left(\frac{k}{\mathcal{H}(t_{e})}\right)^{3}\quad\text{for}\quad k<k_{e},$ (3.15)
where $k_{e}=a/\bar{L}(t_{e})\sim\mathcal{H}(t_{e})$, and $t_{e}$ denotes the horizon crossing time of the wall. The second approximate equality considers a horizon-sized domain wall during the horizon crossing. The power spectrum for the density contrast is then given by
$\mathcal{P}_{\delta}(k)=f_{w}^{2}(t)\mathcal{P}_{\delta_{w}}(k)\simeq f_{w}^{2}(t)\left(\frac{k}{\mathcal{H}(t_{e})}\right)^{3}\quad\text{for}\quad k\lesssim\mathcal{H}(t_{e}).$ (3.16)
As discussed above, the horizon crossing of the domain wall occurs when the pressure from the bias potential becomes comparable to the surface energy of the wall. Therefore, it is expected that the annihilation of the domain wall takes place near the time of the horizon crossing, i.e., $t_{e}\simeq t_{\rm ann}$. We also consider the scenario in which PBHs are formed at the horizon crossing time of the wall. Thus, we have $t_{f}=t_{e}\simeq t_{\rm ann}$. For simplicity, in the following analysis, we adopt the domain wall annihilation time $t_{\rm ann}$ for evaluating the power spectrum.
In the inflationary theory for PBH production, the variance of the density fluctuations is generally related to the power spectrum by
$\sigma^{2}=\int_{0}^{\infty}W^{2}(kR)\,\mathcal{P}_{\delta}(k)\,\frac{dk}{k},$ (3.17)
where $W(kR)$ is a volume-normalized Fourier transform of the window function smoothed over a comoving scale $R$, often assumed to be Gaussian. Instead of using Eq. (3.17), which depends on the assumption of $W(kR)$, we adopt Eq. (3.6) as the variance of the density perturbation from the domain wall network during its horizon crossing
$\sigma^{2}=\left.\langle\delta^{2}\rangle\right|_{V\to V_{H}}=\frac{f_{w}^{2}}{N}.$ (3.18)
Note that if the primordial black hole is formed during the horizon crossing of the domain wall, i.e., $k\simeq\mathcal{H}(t_{e})$, the power spectrum is $\mathcal{P}_{\delta}(k)\simeq f_{w}^{2}(t_{e})$. We observe that the variances given by Eqs. (3.17) and (3.18) are of the same order, $\sim f_{w}^{2}(t_{e})$, except for a statistical factor $1/N$. The case of the horizon crossing of the domain wall and the superhorizon mode $k$ occurring at the same time $t_{e}=t_{f}\simeq t_{\rm ann}$ provides the primary scenario for the production of PBHs, as will be considered below.
Then the probability density distribution function of the density contrast, $P(\delta)$, is featured by a zero ensemble averaged value, $\langle\delta\rangle=0$, and a variance $\sigma$ given by Eq. (3.18). In this work, we consider a Gaussian density distribution function
$P(\delta)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{\delta^{2}}{2\sigma^{2}}\right).$ (3.19)
@@
Simulations [54, 55, 49] indicate that the domain walls tend to merge with the neighboring ones in the evolution, resulting in a wall network with a small number of objects. For a model with a $Z_{2}$ symmetry, there eventually exists $N=1$ domain wall that separates two domains [54]. For models with $\mathcal{N}\sim 10$ domains, one expects that $N\sim\mathcal{N}$ domain walls exist in the late stage of the network evolution. Therefore, once the domain walls contribute significantly to the total energy density, with $f_{w}\sim 0.1$, near the time of domain wall annihilation, the energy density perturbation arising from the number density fluctuations of the domain walls can become remarkable.
4 Isocurvature perturbation from domain wall network
In this section, we consider cosmological perturbations in the background spacetime, described by the flat Friedmann–Robertson–Walker (FRW) Universe. The FRW metric, including the first-order scalar perturbations in the conformal-Newtonian gauge, is given by
$ds^{2}=a(\eta)^{2}\left[-(1+2\Phi)d\eta^{2}+(1-2\Psi)\delta_{ij}dx^{i}dx^{j}\right],$ (4.1)
where $\Phi$ represents the Newtonian potential and $\Psi$ denotes the intrinsic curvature perturbation, both of which are gauge-invariant quantities. The background evolution is governed by the Friedmann equation
$\mathcal{H}^{2}=\frac{8\pi G}{3}a^{2}\bar{\rho},$ (4.2)
and the energy continuity equation
$\bar{\rho}^{\prime}=-3\mathcal{H}\bar{\rho}(1+\omega),$ (4.3)
where $\omega=\bar{p}/\bar{\rho}$ is the equation of state parameter, and the prime denotes the derivative with respect to the conformal time. From these equations, we can derive several useful background relations. Here, we list two of them
$\frac{w^{\prime}}{1+w}$ $=$ $3\mathcal{H}\left(w-c_{s}^{2}\right)$ (4.4)
$\bar{p}^{\prime}=w\bar{\rho}^{\prime}+w^{\prime}\bar{\rho}$ $=$ $-3\mathcal{H}(1+w)c_{s}^{2}\bar{\rho},$ (4.5)
where the sound speed $c_{s}^{2}=\bar{p}^{\prime}/\bar{\rho}^{\prime}$. In Fourier space, the Einstein equations for scalar perturbations are given by [56]
$\left(\frac{k}{\mathcal{H}}\right)^{2}\Psi=-\frac{3}{2}\left[\delta^{N}+3(1+w)\frac{\mathcal{H}}{k}v^{N}\right],$ (4.8)
$\left(\frac{k}{\mathcal{H}}\right)^{2}(\Psi-\Phi)=3w\Pi,$
$\mathcal{H}^{-1}\Psi^{\prime}+\Phi=\frac{3}{2}(1+w)\frac{\mathcal{H}}{k}v^{N},$
$\mathcal{H}^{-2}\Psi^{\prime\prime}+\mathcal{H}^{-1}\left(\Phi^{\prime}+2\Psi^{\prime}\right)+\left(1+\frac{2\mathcal{H}^{\prime}}{\mathcal{H}^{2}}\right)\Phi-\frac{1}{3}\left(\frac{k}{\mathcal{H}}\right)^{2}(\Phi-\Psi)=\frac{3}{2}\frac{\delta p^{N}}{\bar{\rho}},$ (4.9)
where the subscript $N$ denotes the conformal-Newtonian gauge. Here, the relative energy density perturbation is defined as $\delta=\delta\rho/\bar{\rho}$ with $\delta\rho=\rho-\bar{\rho}$, $v$ is the fluid velocity, and $\Pi$ represents the anisotropic stress of the fluid. The energy-momentum continuity equation $T^{\mu}_{\nu;\mu}=0$ yields the density perturbations in the conformal-Newtonian gauge
$\left(\delta^{N}\right)^{\prime}$ $=$ $(1+w)\left(-kv^{N}+3\Psi^{\prime}\right)+3\mathcal{H}\left(w\delta^{N}-\frac{\delta p^{N}}{\bar{\rho}}\right),$ (4.10)
$\left(v^{N}\right)^{\prime}$ $=$ $-\mathcal{H}(1-3w)v^{N}-\frac{w^{\prime}}{1+w}v^{N}+\frac{k\delta p^{N}}{\bar{\rho}+\bar{p}}-\frac{2}{3}\frac{w}{1+w}k\Pi+k\Phi.$ (4.11)
Let us consider a two-component universe with the total fluid energy density and pressure given by $\bar{\rho}=\sum\bar{\rho}_{i}$ and $\bar{p}=\sum\bar{p}_{i}$, respectively. The component fluids are labeled by $i$ and $j$, while quantities without a subscript denote variables of the total fluid. We assume that energy transfer between the fluid components in the background universe is negligible. In this case, each component fluid separately satisfies the energy continuity equation (4.3), the background relations (4.4)-(4.5), and the energy-momentum continuity equations (4.10)-(4.11). The quantities for the total and component fluids are related by
$\omega=\frac{\sum\bar{\rho}_{i}\omega_{i}}{\bar{\rho}},\leavevmode\nobreak\ \leavevmode\nobreak\ c_{s}^{2}=\sum\frac{\bar{\rho}_{i}+\bar{p}_{i}}{\bar{\rho}+\bar{p}}c_{i}^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \delta=\frac{\sum\bar{\rho}_{i}\delta_{i}}{\bar{\rho}},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ v=\sum\frac{1+w_{i}}{1+w}\frac{\bar{\rho}_{i}}{\bar{\rho}}v_{i}.$ (4.12)
We also define the total entropy perturbation as
$\mathcal{S}\equiv\mathcal{H}\left(\frac{\delta p}{\bar{p}^{\prime}}-\frac{\delta\rho}{\bar{\rho}^{\prime}}\right)=\frac{1}{3(1+w)}\left(\frac{\delta\rho}{\bar{\rho}}-\frac{1}{c_{s}^{2}}\frac{\delta p}{\bar{\rho}}\right).$ (4.13)
Here, we have used Eq. (4.3) and $\bar{p}^{\prime}=c_{s}^{2}\bar{\rho}$ for the second equality. This definition is also valid for the component fluids, as there is no energy transfer between them. The relative entropy perturbation, or isocurvature perturbation, is defined as
$S_{ij}\equiv\mathcal{H}\left(\frac{\delta\rho_{i}}{\bar{\rho}_{i}^{\prime}}-\frac{\delta\rho_{j}}{\bar{\rho}_{j}^{\prime}}\right)=\frac{\delta_{i}}{1+\omega_{i}}-\frac{\delta_{j}}{1+\omega_{j}}.$ (4.14)
In the following, we omit the subscript and use $S$ to denote the isocurvature perturbation.
Using Eqs. (4.4), (4.10), and (4.13), we obtain
$S^{\prime}=-kv_{\rm rel}-9\mathcal{H}\left(c_{i}^{2}\mathcal{S}_{i}-c_{j}^{2}\mathcal{S}_{j}\right),$ (4.15)
where the relative velocity $v_{\rm rel}\equiv v_{i}^{N}-v_{j}^{N}=v_{i}-v_{j}$ is a gauge-independent quantity. We further assume the relation $p_{i}=\omega_{i}\bar{\rho}_{i}=c_{i}^{2}\bar{\rho}_{i}$ for both fluid components. Consequently, the internal entropy perturbation $\mathcal{S}_{i}$ in Eq. (4.15) vanishes. Using Eqs. (4.11) and (4.15), we derive the second derivative of $S$
$S^{\prime\prime}=-kv_{\rm rel}^{\prime}=$ $\leavevmode\nobreak\ k\mathcal{H}\left[\left(1-3w_{i}\right)v_{i}^{N}-\left(1-3w_{j}\right)v_{j}^{N}\right]+k\left(\frac{w_{i}^{\prime}}{1+w_{i}}v_{i}^{N}-\frac{w_{j}^{\prime}}{1+w_{j}}v_{j}^{N}\right)$ (4.16)
$-k^{2}\left(\frac{c_{i}^{2}\delta_{i}^{N}}{1+w_{i}}-\frac{c_{j}^{2}\delta_{j}^{N}}{1+w_{j}}\right)+\frac{2}{3}k^{2}\left(\frac{w_{i}}{1+w_{i}}\Pi_{i}-\frac{w_{j}}{1+w_{j}}\Pi_{j}\right).$
In the above derivation, the momentum transfer between components has been neglected. Defining the energy fraction of component $i$ as $f=\bar{\rho}_{i}/\bar{\rho}_{j}$, we can express the component perturbations in terms of the total and relative perturbations through the following transformations
$\left\{\begin{matrix}\delta_{i}=a_{11}\delta+a_{12}S,\\ \delta_{j}=a_{21}\delta+a_{22}S,\end{matrix}\right.\quad\quad{\rm and}\quad\quad\left\{\begin{matrix}v_{i}=b_{11}v+b_{12}v_{\rm rel},\\ v_{j}=b_{21}v+b_{22}v_{\rm rel},\end{matrix}\right$. (4.17)
where
$\left\{\begin{matrix}a_{11}=\frac{(1+\omega_{i})(1+f)}{\lambda},\leavevmode\nobreak\ \\ a_{12}=\frac{(1+\omega_{i})(1+\omega_{j})}{\lambda},\end{matrix}\right.\quad\quad{\rm and}\quad\quad\left\{\begin{matrix}a_{21}=-\frac{(1+\omega_{j})(1+f)}{\lambda},\\ a_{22}=\frac{(1+\omega_{i})(1+\omega_{j})f}{\lambda},\end{matrix}\right$. (4.18)
and
$\left\{\begin{matrix}b_{11}=\frac{(1+\omega)(1+f)}{\lambda},\\ b_{12}=\frac{1+\omega_{j}}{\lambda},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \end{matrix}\right.\quad\quad{\rm and}\quad\quad\left\{\begin{matrix}b_{21}=\frac{(1+\omega)(1+f)}{\lambda},\\ b_{22}=\frac{(1+\omega_{i})f}{\lambda},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \end{matrix}\right$. (4.19)
with $\lambda=1+f+f\omega_{i}+\omega_{j}$. We have used the relation $c_{i}^{2}=\omega_{i}$, which implies $\omega_{i}^{\prime}/(1+\omega_{i})=0$. Neglecting the anisotropic stress and replacing the component quantities with the total and relative ones in Eq. (4.16), we obtain
$\frac{S^{\prime\prime}}{\mathcal{H}^{2}}+\frac{(1-3\mu_{2})S^{\prime}}{\mathcal{H}}=-\left(\frac{k}{\mathcal{H}}\right)^{2}\left(\mu_{1}\delta^{C}+\mu_{2}S\right),$ (4.20)
where
$\delta^{C}=\delta^{N}+3\left(\frac{\mathcal{H}}{k}\right)(1+\omega)v^{N},$ (4.21)
with $C$ representing the comoving gauge, and the parameters
$\mu_{1}=\frac{(\omega_{i}-\omega_{j})(1+f)}{\lambda}\quad{\rm and}\quad\mu_{2}=\frac{\omega_{i}(1+\omega_{j})-\omega_{j}(1+\omega_{i})f}{\lambda}.$ (4.22)
If the components are cold matter and radiation, the parameters reduce to $\mu_{1}=-1/(1+w)$ and $\mu_{2}=(1-3c_{s}^{2})/3$. In this case, we reproduce the so-called Kodama-Sasaki equation [57].
The Bardeen equation determines the evolution of $\delta^{C}$ in the comoving gauge
$\mathcal{H}^{-2}\delta^{C\prime\prime}+\left(1-6w+3c_{s}^{2}\right)\mathcal{H}^{-1}\delta^{C\prime}-\frac{3}{2}\left(1+8w-6c_{s}^{2}-3w^{2}\right)\delta^{C}$ (4.23)
$=-c_{s}^{2}\left(\frac{k}{\mathcal{H}}\right)^{2}\left[\delta^{C}-3(1+w)\mathcal{S}\right].$
With Eqs. (4.17)-(4.19) and considering the conditions $f\ll 1$ and $\delta^{C}\ll S$ for the isocurvature perturbation at the initial time, the Bardeen equation can be written as
$\mathcal{H}^{-2}\delta^{C\prime\prime}-\frac{3\delta^{C\prime}}{4\mathcal{H}}-\frac{17}{4}\delta^{C}=-\frac{1}{12}\left(\frac{k}{\mathcal{H}}\right)^{2}\left(\delta^{C}-\frac{14}{3}fS\right),$ (4.24)
where we have assumed $\omega_{w}=-2/3$ for the domain wall. Together with Eq. (4.20), we can obtain the solution for the evolution of $\delta^{C}$ and $S$. Let us consider the isocurvature modes, i.e., $\delta^{C}\ll S$ and the condition $f\to 0$. Then Eqs. (4.24) and (4.20) can be reduced to
$\mathcal{H}^{-2}\delta^{C\prime\prime}-\frac{3\delta^{C\prime}}{4\mathcal{H}}-\frac{17}{4}\delta^{C}$ $=$ $\frac{17}{48}\left(\frac{k}{\mathcal{H}}\right)^{2}fS,$ (4.25)
$\mathcal{H}^{-2}S^{\prime\prime}+3\mathcal{H}^{-1}S^{\prime}$ $=$ $0.$ (4.26)
Requiring $S(0)$ to be finite at the initial time $\tau=0$, the solution to Eq. (4.26) is a constant, i.e., $S=S(0)$. With this result, we can determine $\delta^{C}$ by solving Eq. (4.25), yielding
$\delta^{C}(\tau)=\frac{7}{90}k^{2}\tau^{2}f_{w}S(0).$ (4.27)
Using the Einstein equations (4.8)-(4.9), we derive various relations among perturbation quantities, summarized as follows
$\Phi$ $=$ $-\frac{3}{2}\left(\frac{\mathcal{H}}{k}\right)^{2}\delta^{C},$ (4.28)
$v^{N}$ $=$ $\frac{2}{3(1+w)}\left(\frac{k}{\mathcal{H}}\right)\left(\mathcal{H}^{-1}\Phi^{\prime}+\Phi\right),$ (4.29)
$\delta^{N}$ $=$ $\delta^{C}-3\left(\frac{\mathcal{H}}{k}\right)(1+w)v^{N},$ (4.30)
where we have assumed a radiation-dominated universe, in which the anisotropic stress is negligible and $\Phi=\Psi$. Together with Eq. (4.27), we find
$\delta^{N}\equiv\delta\simeq\frac{1}{3}f_{w}S(0)\quad{\rm for}\quad k\to 0.$ (4.31)
For the superhorizon mode of the isocurvature perturbation $k\ll\mathcal{H}(t)$, we can neglect the term on the right-hand side of Eq. (4.20). Considering a radiation-dominated universe and requiring $S(0)$ to be finite with $S^{\prime}(0)=0$, we again find the solution $S=S(0)$ for the superhorizon mode.
We assume that the density perturbation is primarily induced by the Poisson fluctuation of the domain wall network and neglect the radiation density fluctuation. Consequently, $S\simeq\delta_{w}/(1+\omega_{w})=3\delta/f_{w}$. Based on the discussions in Section 2.2, we find that the domain walls are initially generated as a radiation gas, and therefore, the isocurvature perturbations are not produced at the initial time. The size (correlation length) of the domain wall rapidly grows to a superhorizon scale, and the closed configurations are formed before the domain wall is slowed down. The correlation length of the wall can be significantly reduced when the bias potential plays an important role in the domain wall evolution. As a result, isocurvature perturbations are produced only during the horizon crossing of the domain wall, which occurs near the annihilation time of the domain wall, as the bias potential has already become comparable to the surface tension of the wall. This implies that the superhorizon mode of the density perturbation arising from the isocurvature perturbation of the wall is given by
$\delta=f_{w}(t_{\rm ann})\delta_{w}\simeq\frac{1}{3}f_{w}(t_{\rm ann})S(0),$ (4.32)
where the relation $t_{e}=t_{f}\simeq t_{\rm ann}$ has been adopted. With $\Phi=\Psi$ in the radiation-dominated universe, Eq. (4.8) provides the Newtonian potential during the horizon crossing of mode $k$
$\Phi(t_{f})\simeq-\frac{3}{2}\left(\frac{k}{\mathcal{H}(t_{\rm ann})}\right)^{2}\delta^{N}(t_{\rm ann}).$ (4.33)
Furthermore, from Eq. (4.21), we also have $\delta^{C}\sim\delta^{N}$ when $k\sim\mathcal{H}(t_{\rm ann})$. Therefore, we expect that the gauge dependence of the density contrast becomes weak during the horizon crossing of mode $k$.